The cd's at a store were on sale for $10 and $12. THE counter at the exit says 500 people left with a cd (limit 1 per customer) and the register contains $5750. Maria wrote a system of equations for the number of $10 and $12 CD's. What was the system of equations.
Answer by checkley75(3666)

X+Y=500 X=500-Y
10X+12Y=5750 NOW REPLACE X WITH (500-Y) & SOLVE FOR Y
10(500-Y)+12Y=5750
5000-10Y+12Y=5750
2Y=5750-5000
2Y=750
Y=750/2
Y=375 THE NUMBER OF $12 CD'S.
NOW SUBSTITUTE 12 FOR Y & SOLVE FOR X
X=275=500
X=500-275
X=225 THE NUMBER OF $10 CD'S.
PROOF
10*225+12*275=5750
2250+3300=5750
5750=5750